# Adaptive Visual Scene Understanding: Incremental Scene Graph Generation

### v1

@@ A.6.2 Learning Scenario 2 (S2)
From Fig.S19 we observe that, in Task 1, the ground truth scene graph contains triplets: “man riding skateboard”, “man above skateboard”, and “shoe of skateboard”. After training on task 1 data, all three models (Replay 10%, Naive, EWC) predict “man riding skateboard” and “man above skateboard”. None of the models predict “shoe of skateboard”.
@@
In Task 2, new triplets introduced are: “leg of person”, “person wearing sock”, “person has head”, and “head of person”. After training on task 2 data, the Replay 10% model only remembers “man riding skateboard” from the previous task, forgetting “man above skateboard”. Moreover, Replay 10% model can only predict “leg of person” from the triplets introduced in task 2. The Naive model forgets all the triplets from task 1 ( “man riding skateboard”, “man above skateboard”) and only predicts “leg of person” from the triplets introduced in task 2. Similar to Naive model, EWC model forgets all the triplets from task 1 ( “man riding skateboard”, “man above skateboard”) and only predicts “leg of person” from the triplets introduced in task 2.
@@
Figure S20: Visualization example for Learning Scenario 3 (S3). The leftmost column shows the standalone ground truth bounding boxes and scene graphs in the generalization test set regardless of which task it is in Learning Scenario 3 (S3). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Naive, Replay 10%). Similar to Fig. S18, and S19, the red boxes in all scene graphs indicate objects , while green boxes represent relationships. As explained in Fig. S18 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. The time arrow on the left side of the figure demonstrates that the model is exposed to new objects over time as described in Sec. 3.1. For easy referral of object instances in the predicted scene graphs, we numbered the unknown bounding boxes in this figure, where the numbers are not actually present in the model predictions. Concretely, unk_1 refers to "mountain"; unk_2 is "house"; and unk_3 is "sheep".
A.6.3 Learning Scenario 3 (S3)
@@
Fig. S20 illustrates the performance of the Replay 10% and Naive models in locating unknown objects and recognizing the relationships between these objects and other nearby unknown objects. The ground truth in Fig. S20 consists of three unknown objects: "mountain", "sheep", and "house", along with three relationships: "near", "behind", and "infront of" (between "mountain" and "house"), and a "near" relationship (between "sheep" and "house").
@@
After training on Task 1 data, the Naive model can accurately locate the objects "mountain" and "house". However, no new object, such as "sheep", is located by the Naive model after training on Task 2 data. After training on Task 3 data, the Naive model is also able to locate the object "sheep" in addition to "mountain" and "house". Even after training on Task 4 data, the Naive model continues to locate all three objects: "mountain", "house", and "sheep". In contrast, the Replay 10% model, after training on Task 1 data, can only locate "mountain" and "house". This remains the same even after training on Task 2 and Task 3 data, where the Replay 10% model can still only locate the objects "mountain" and "house". However, after training on Task 4 data, the Replay 10% model is able to locate all the objects: "mountain", "house", and "sheep".
@@
Regarding relationship generalization on unknown objects, the Naive model, after training on Task 1, can only predict the "near" relationship between the located objects "mountain" and "house" out of the three possible relationships. This performance remains the same even after Task 2. However, after training on Task 3, the Naive model can predict the "behind" relationship in addition to the "near" relationship between the located objects "mountain" and "house". After Task 4, the Naive model can predict the "behind" and "near" relationships between the located objects "mountain" and "house", as well as the "near" relationship between the located objects "sheep" and "house". In contrast, the Replay 10% model, after training on Task 1, can predict the "near" and “behind” relationships between the located objects "mountain" and "house". After Task 2, it can also predict the "infront of" relationship between the located objects "mountain" and "house" along with “near” and “behind” relationships. Even after Task 3, the Replay 10% model is still able to predict the "near", "behind", and "infront of" relationships between the located objects "mountain" and "house". After Task 4, as it can now locate the object "sheep", the Replay 10% model can also predict the "near" relationship between the objects "house" and "sheep", in addition to the existing relationships between "mountain" and "house".

### v3

@@ 0.A.5.2 Learning Scenario 2 (S2)
From Fig.S6, we observe that, in Task 1, the ground truth scene graph contains triplets: “man riding skateboard”, “man above skateboard”, and “shoe of skateboard”. After training on task 1 data, all three models (Replay@10%, Naive, EWC) predict “man riding skateboard” and “man above skateboard”. None of the models predict “shoe of skateboard”.
@@
In Task 2, new triplets introduced are: “leg of person”, “person wearing sock”, “person has head”, and “head of person”. After training on task 2 data, the Replay@10% model only remembers “man riding skateboard” from the previous task, forgetting “man above skateboard”. Moreover, Replay@10% model can only predict “leg of person” from the triplets introduced in task 2. The Naive model forgets all the triplets from task 1 ( “man riding skateboard”, “man above skateboard”) and only predicts “leg of person” from the triplets introduced in task 2. Similar to Naive model, EWC model forgets all the triplets from task 1 ( “man riding skateboard”, “man above skateboard”) and only predicts “leg of person” from the triplets introduced in task 2.
@@
Figure S7: Visualization example for Learning Scenario 3 (S3). The leftmost column shows the standalone ground truth bounding boxes and scene graphs in the generalization test set regardless of which task it is in Learning Scenario 3 (S3). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Naive, Replay@10%). Similar to Fig. S5, and S6, the red boxes in all scene graphs indicate objects , while green boxes represent relationships. As explained in Fig. S5 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. The time arrow on the left side of the figure demonstrates that the model is exposed to new objects over time as described in Sec. 3.1. For easy referral of object instances in the predicted scene graphs, we numbered the unknown bounding boxes in this figure, where the numbers are not actually present in the model predictions. Concretely, unk_1 refers to "mountain"; unk_2 is "house"; and unk_3 is "sheep".
0.A.5.3 Learning Scenario 3 (S3)
@@
Fig. S7 illustrates the performance of the Replay@10% and Naive models in locating unknown objects and recognizing the relationships between these objects and other nearby unknown objects. The ground truth in Fig. S7 consists of three unknown objects: "mountain", "sheep", and "house", along with three relationships: "near", "behind", and "infront of" (between "mountain" and "house"), and a "near" relationship (between "sheep" and "house").
@@
After training on Task 1 data, the Naive model can accurately locate the objects "mountain" and "house". However, no new object, such as "sheep", is located by the Naive model after training on Task 2 data. After training on Task 3 data, the Naive model is also able to locate the object "sheep" in addition to "mountain" and "house". Even after training on Task 4 data, the Naive model continues to locate all three objects: "mountain", "house", and "sheep". In contrast, the Replay@10% model, after training on Task 1 data, can only locate "mountain" and "house". This remains the same even after training on Task 2 and Task 3 data, where the Replay@10% model can still only locate the objects "mountain" and "house". However, after training on Task 4 data, the Replay@10% model is able to locate all the objects: "mountain", "house", and "sheep".
@@
Regarding relationship generalization on unknown objects, the Naive model, after training on Task 1, can only predict the "near" relationship between the located objects "mountain" and "house" out of the three possible relationships. This performance remains the same even after Task 2. However, after training on Task 3, the Naive model can predict the "behind" relationship in addition to the "near" relationship between the located objects "mountain" and "house". After Task 4, the Naive model can predict the "behind" and "near" relationships between the located objects "mountain" and "house", as well as the "near" relationship between the located objects "sheep" and "house". In contrast, the Replay@10% model, after training on Task 1, can predict the "near" and “behind” relationships between the located objects "mountain" and "house". After Task 2, it can also predict the "infront of" relationship between the located objects "mountain" and "house" along with “near” and “behind” relationships. Even after Task 3, the Replay@10% model is still able to predict the "near", "behind", and "infront of" relationships between the located objects "mountain" and "house". After Task 4, as it can now locate the object "sheep", the Replay@10% model can also predict the "near" relationship between the objects "house" and "sheep", in addition to the existing relationships between "mountain" and "house".
0.A.6 Ethical Concerns
The development and deployment of Scene Graph Generation (SGG) technology present potential negative societal impacts that warrant careful consideration [28]. Firstly, privacy concerns arise as SGG may inadvertently capture sensitive information from images, potentially violating privacy rights and raising surveillance issues. Secondly, bias and fairness challenges persist, as SGG algorithms can perpetuate biases present in training data, leading to discriminatory outcomes that reinforce societal inequalities. Misinterpretation and misclassification by SGG algorithms could result in misinformation and incorrect actions, impacting decision-making. The risk of manipulation and misuse of SGG-generated scene representations for malicious purposes is also a concern. For example, attackers might manipulate scene graphs to deceive systems or disrupt applications that rely on scene understanding.